OpenStudy (anonymous):

5(2x - 7) = 33 - 3(6x + 4) show all work! distribute and solve for x

OpenStudy (anonymous):

5(2X-7)=33-3(6X+4) 10X-35=33-18X-12 10X-35=21-18X +18X +18X 28X-35=21 +35 +35 28X=56 28/28 56/28 X=2